OpenStudy (anonymous):

Find the distance between the points (-4, 6) and (8, -12).Round your solution to 2 decimal points.

OpenStudy (anonymous):

you can get a ruler to measure it

OpenStudy (anonymous):

i dont have a ruler

OpenStudy (jdoe0001):

\(\large \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-4\quad ,&6)\quad &(8\quad ,&-12) \end{array}\qquad d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2} \)

OpenStudy (anonymous):

It only cost like $1.99 at walmart

OpenStudy (anonymous):

I am aware of how much one is.

OpenStudy (jdoe0001):

$1.99 for a ruler? I can find it cheaper than that for $1.99 they must make some profit I'd think, I mean, they paid about 25cents

OpenStudy (texaschic101):

d = sq rt (x2 - x1)^2 + (y2 - y1)^2 d = sq rt (8 - (-4)^2 + (-12 - 6)^2 d = sq rt (8 + 4)^2 + (-18^2) d = sq rt 144 + 324 d = sq rt 468 d = 21.633 rounds to 21.63